OpenStudy (henrietepurina):

Part 1. Create two radical equations: one that has an extraneous solution, and one that does not have an extraneous solution. Use the equation below as a model. a(√x+b)+c=d Use a constant in place of each variable a, b, c, and d. You can use positive and negative constants in your equation. Part 2. Show your work in solving the equation. Include the work to check your solution and show that your solution is extraneous. Part 3. Explain why the first equation has an extraneous solution and the second does not.

hero (hero):

I can help you tomorrow. I have to go now. Sorry.

OpenStudy (henrietepurina):

@SolomonZelman

OpenStudy (henrietepurina):

Can you help?

OpenStudy (henrietepurina):

I really need this for tommorow...

OpenStudy (henrietepurina):

hello? @SolomonZelman

OpenStudy (solomonzelman):

there is no reason to tag me. Sorry.

OpenStudy (henrietepurina):

ok im sorry...

OpenStudy (henrietepurina):

can we start with 1?

OpenStudy (henrietepurina):

@SolomonZelman

OpenStudy (solomonzelman):

for part one, what have you tried to do ?

OpenStudy (henrietepurina):

ok... i wrote 1 equation: 2(√x+3)+4=5 That is a extraenous one but i cant think of one thats not

OpenStudy (solomonzelman):

yeah, kind of hard to do that... anyway , for part 2, what do you get ?

OpenStudy (solomonzelman):

I would write 2(√x+3)+1=9

OpenStudy (henrietepurina):

for non extraneous?

OpenStudy (solomonzelman):

this is for extraneous. 2(√x+3)+1=9 solution, 2(√x+3)=9-1 2(√x+3)=8 (√x+3)=4 x+3=16 x=13

OpenStudy (henrietepurina):

ok thanks but what about the steps for the extraneous one? 2(√x+3)+4=5 2(√x+3)=1 √x+3=0.5 x+3=0.5 x=-3.5?

OpenStudy (solomonzelman):

I AM thinking... 2(√x+1)+3=-5 2(√x+1)=-5-3 2(√x+1)=-8 (√x+1)=-4

OpenStudy (solomonzelman):

and no x-value will make the root be equal to -4 :)

OpenStudy (henrietepurina):

as extraneous?

OpenStudy (solomonzelman):

the last one?

OpenStudy (henrietepurina):

yup

OpenStudy (solomonzelman):

Well if you have 2(√x+1)+3=-5 with the solution below 2(√x+1)=-5-3 2(√x+1)=-8 (√x+1)=-4 and knowing that such a value that makes the root =-4 doesn't exist what do you think ?

OpenStudy (solomonzelman):

no hold on, sorry

OpenStudy (solomonzelman):

it does work if x=15

OpenStudy (solomonzelman):

b/c √16=±4

OpenStudy (solomonzelman):

idk the non extraneous one.

OpenStudy (henrietepurina):

hmm... what about well we already had one remember non extraneous means a solution and that one had it :)

OpenStudy (solomonzelman):

sorry, I can't help . I did whatever I can.